Buzz Appliances manufactures two​ products: Food Processors and Espresso Machines. The following data are​ available:

Food Processors

Espresso Makers

Sales price

​$125

​$265

Variable costs

​$50

​$170

The company can manufacture either two food processors per machine hour or three espresso machines per machine hour. The​ company's production capacity is​ 1,400 machine hours per month.

The company has demand of​ 2,700 espresso machines. How many espresso machines and food processors should they produce based on demand and available machine​ hours?

A.

900 espresso machines and 500 food processors

B.

​2,700 espresso machines and 500 food processors

C.

​2,700 espresso machines and 0 food processors

D.

​2,700 espresso machines and​ 1,000 food processors
Answer #1

Food Processors

Espresso Makers

Sales price

​$125

​$265

Variable costs

​$50

​$170

Contribution margin per unit $75 $95
Machine hour per unit 0.5 1/3
Contribution margin per Machine hour 75/0.5 = $150 95 x 3 = $285

Since Contribution margin per Machine hour is higher for Espresso Makers, hence Espresso Makers will be made to the maximum possible extent and from remaining machine hours, Food Processors will be made.

The company has demand of​ 2,700 espresso machines

Machine hours used in making 2,700 espresso machines = 2,700 x 1/3

= 900 hours

Remaining machine hours = 1,400 - 900

= 500 hours

Production of Food Processors = 500/0.5

= 1,000

Hence, correct option is (D) i.e. 2,700 espresso machines and​ 1,000 food processors

Please ask if you have any query related to the question. Thank you